# Continual Learning: Forget-free Winning Subnetworks for Video Representations

### v3

@@ Abstract
Inspired by the Lottery Ticket Hypothesis (LTH), which highlights the existence of efficient subnetworks within larger, dense networks, a high-performing Winning Subnetwork (WSN) in terms of task performance under appropriate sparsity conditions is considered for various continual learning tasks. It leverages pre-existing weights from dense networks to achieve efficient learning in Task Incremental Learning (TIL) scenarios. In Few-Shot Class Incremental Learning (FSCIL), a variation of WSN referred to as the Soft subnetwork (SoftNet) is designed to prevent overfitting when the data samples are scarce. Furthermore, the sparse reuse of WSN weights is considered for Video Incremental Learning (VIL). The use of Fourier Subneural Operator (FSO) within WSN is considered. It enables compact encoding of videos and identifies reusable subnetworks across varying bandwidths. We have integrated FSO into different architectural frameworks for continual learning, including VIL, TIL, and FSCIL. Our comprehensive experiments demonstrate FSO’s effectiveness, significantly improving task performance at various convolutional representational levels. Specifically, FSO enhances higher-layer performance in TIL and FSCIL and lower-layer performance in VIL.
@@
Continual Learning (CL), Task Incremental Learning (TIL), Video Incremental Learning (VIL), Few-shot Class Incremental Learning (FSCIL), Regularized Lottery Ticket Hypothesis (RLTH), Wining SubNetworks (WSN), Soft-Subnetwork (SoftNet), Fourier Subneural Operator (FSO)
@@ 1 Introduction
These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting or to alleviate overfitting. However, they are unsuited for sequential complex Video Incremental Learning (VIL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complex signals) in frequency space. To achieve this, we define Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. Given a backbone and FSO architecture, our method continuously learns to identify input-adaptive sub-modules in Fourier space and encode videos in each sequential training session. We apply FSO to various architectures accompanied by continual learning scenarios, such as Task Incremental Learning (TIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO representations.
@@
  • We have applied the FSO to various architectures used by a variety of continual learning scenarios: Video Incremental Learning (VIL), Task Incremental Learning (TIL), and Few-Shot Class Incremental Learning (FSCIL). In our evaluations, the proposed FSO performs better than architecture-based continual learning models, such as WSN and SoftNet, in TIL, VIL, and FSCIL, respectively, underscoring its exceptional representational power.
@@ 2 Related Works
Continual Learning [56, 83, 42, 46], also known as lifelong learning, is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is regularization-based methods [38, 4, 31, 84, 60], which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. Rehearsal-based approaches [65, 5, 6, 68, 17, 80, 69] store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some methods in this line of work [74, 2] accommodate the generative model to construct the pseudo-rehearsals for previous tasks. Architecture-based approaches [53, 71, 45, 90, 34, 35] use the additional capacity to expand [92, 97], dynamic representation [93, 75] or isolate [67] model parameters, preserving learned knowledge and preventing forgetting. Both rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters [90] or the replay buffers. Recently, Prompt-based learning, an emerging transfer learning technique in natural language processing (NLP), harnesses a fixed function of pre-trained Transformer models. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Notably, while L2P [89] stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt [88] introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox [19], S-Prompt [87], CODA-P [78], ConStruct-VL [77], ST-Prompt [62], and LGCL [36]. All previous learning methods depend mainly on continual representations in real space. However, the central focus of this study is to pinpoint the most optimal winning ticket representations for convolutional operators in three continual learning scenarios in Fourier space.
Architecture-based Continual Learning. While most works aim to increase the performance of continual learners by adding memory, some researchers have focused on building memory and computationally efficient continual learners by using pruning-based constraints. CLNP [21] is one example of a method that selects important neurons for a given task using $\ell_{1}$ regularization to induce sparsity and freezes them to maintain performance. Neurons that are not selected are reinitialized for future task training. Another method, Piggyback [53], trains task-specific binary masks on the weights given a pre-trained model. However, this method does not allow for knowledge transfer among tasks, and its performance highly depends on the quality of the backbone model. HAT [71] proposes task-specific learnable attention vectors to identify significant weights per task. The masks are formulated to layerwise cumulative attention vectors during continual learning. A recent method, LL-Tickets [12], shows a sparse subnetwork called lifelong tickets that performs well on all tasks during continual learning. The method searches for more prominent tickets from current ones if the obtained tickets cannot sufficiently learn the new task while maintaining performance on past tasks. However, LL-Tickets require external data to maximize knowledge distillation with learned models for prior tasks, and the ticket expansion process involves retraining and pruning steps. WSN [34] is a new method that jointly learns the model and task-adaptive binary masks on subnetworks associated with each task, which can be extended to Fourier spaces. It attempts to select a small set of weights activated (winning ticket) by reusing weights of the prior subnetworks in Task Incremental Learning (TIL). Also, in Few-Shot Class Incremental Learning (FSCIL), we investigate the performance of Soft-Subnetworks (SoftNet) [35] as another variant of WSN: we select sub-networks (winning tickets) to obtain base session knowledge and optimize the regularized sub-network parameters to alleviate overfitting few samples for new sessions.
@@
Neural Implicit Representation (NIR) [57] are neural network architectures for parameterizing continuous, differentiable signals. Based on coordinate information, they provide a way to represent complex, high-dimensional data with a small set of learnable parameters that can be used for various tasks such as image reconstruction [76, 81], shape regression [13, 61], and 3D view synthesis  [59, 70]. Instead of using coordinate-based methods, NeRV [10] proposes an image-wise implicit representation that takes frame indices as inputs, enabling fast and accurate video compression. NeRV has inspired further improvements in video regression by CNeRV [8], DNeRV [25], E-NeRV [47], and NIRVANA [52], and HNeRV [9]. A few recent works have explored video continual learning (VCL) scenarios for the NIR. To tackle non-physical environments, Continual Predictive Learning (CPL) [7] learns a mixture world model via predictive experience replay and performs test-time adaptation using non-parametric task inference. PIVOT [86] leverages the past knowledge present in pre-trained models from the image domain to reduce the number of trainable parameters and mitigate forgetting. CPL needs memory to replay, while PIVOT needs pre-training and fine-tuning steps. In contrast, along with the conventional progressive training techniques [67, 16], we use WSN as baselines, which utilizes the Lottery Ticket Hypothesis (LTH) to identify an adaptive substructure within the dense networks that are tailored to the specific video input index. However, WSN is inappropriate for sequential complex video compilation since it reuses a few adaptive but sparse learnable parameters. To overcome the weakness of WSN, We proposed a novel Fourier Subneural Operator (FSO) [33] for representing complex video in Fourier space [48, 49, 39, 85]. We have expanded the FSO of Fourier representations to encompass a variety of continual learning architectures and scenarios to validate its effectivness.
@@ 3 Winning SubNetworks in Fourier Space
This section presents our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[34] and introduces the Fourier Subneural Operator (FSO, see Figure 3 and Figure 4) for better video representations. Then, we depict how we apply FSO to various architectures used in three continual learning scenarios.
3.1 WSN & Fourier Subneural Operator (FSO)
The WSN searches for the task-adaptive winning tickets and updates only the weights not trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [35], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting. However, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, VIL, and FSCIL, to demonstrate the effectiveness of FSO.
@@ 3.1 WSN & Fourier Subneural Operator (FSO)
Figure 2: An illustration of Winning SubNetworks (WSN): (a) The top-cweights $\hat{\bm{\theta}}_{s-1}$ at prior task are obtained, (b) In the forward pass of a new task, WSN reuses weights selected from prior tasks, (c) In the backward pass, WSN updates only non-used weights, and (d) after several iterations of (b) and (c), we acquire again the top-cweights $\hat{\bm{\theta}}_{s}$ including subsets of reused weights (green) for the new task.
Problem Statement. Consider a supervised learning setup where $S$ sessions or tasks arrive to a learner sequentially. We denote that $\mathcal{D}_{s}=\{\textbf{x}_{i,s},y_{i,s}\}_{i=1}^{n_{s}}$ is the dataset of session $s$, composed of $n_{s}$ pairs of raw instances and corresponding labels. We assume a neural network $f(\cdot;\bm{\theta})$, parameterized by the model weights $\bm{\theta}$ and standard continual learning scenario aims to learn a sequence of sessions by solving the following optimization procedure at each step $s$:
$\bm{\theta}^{\ast}=\operatorname*{minimize}_{\bm{\theta}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}(f(\textbf{x}_{i,s};\bm{\theta}),y_{i,s}),$ (1)
@@
Continual learners frequently use over-parameterized deep neural networks to ensure enough capacity for learning future tasks. This approach often leads to the discovery of subnetworks that perform as well as or better than the dense network. Given the neural network parameters $\bm{\theta}$, the binary attention mask $\bm{m}^{*}_{s}$ that describes the optimal subnetwork for session $s$ such that $|\bm{m}^{*}_{s}|$ is less than the model capacity $c$ follows as:
$\begin{split}\bm{m}^{*}_{s}&=\underset{\bm{m}_{s}\in\{0,1\}^{|\bm{\theta}|}}{\operatorname*{minimize}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}\big{(}f(\textbf{x}_{i,s};\bm{\theta}\odot\bm{m}_{s}),y_{i,s}\big{)}-\mathcal{J}\\ &\quad\quad\quad\quad\quad\quad\text{subject to\leavevmode\nobreak\ }|\bm{m}^{*}_{s}|\leq c,\end{split}$ (2)
@@ 3.2 Fourier Subnueral Operator (FSO)
$\left(\mathcal{K}(\phi)\bm{e}_{s,t}\right)(\bm{v}_{t}^{s})=\mathcal{F}^{-1}(R_{\bm{\phi}}\cdot(\mathcal{F}\bm{e}_{s,t}))(\bm{v}_{t}^{s}),$ (3)
where $R_{\bm{\phi}}$ is the Fourier transform of a periodic subnetwork function which is parameterized by its subnetwork’s parameters of real $(\bm{\theta}^{real}\odot\bm{m}_{s}^{real})$ and imaginary $(\bm{\theta}^{imag}\odot\bm{m}_{s}^{imag})$. We thus parameterize $R_{\bm{\phi}}$ separately as complex-valued tensors of real and imaginary $\bm{\phi}_{FSO}\in\{\bm{\theta}^{real},\bm{\theta}^{imag}\}$. One key aspect of the FSO is that its parameters grow with the depth of the layer and the input/output size. However, through careful layer-wise inspection and adjustments for sparsity, we can find a balance that allows the FSO to describe neural implicit representations efficiently. In the experimental section, we will showcase the most efficient FSO structure and its performance. Figure 3 shows one possible structure of a single FSO. We describe the optimization in the following section.
Figure 3: Forget-free Neural Implicit Representation with Fourier Subneural Operator (FSO) for Video Incremental Learning: Image-wise neural implicit representation taking frame and video (session $s$) indices as input and using a sparse Stems + NeRV Blocks with Fourier Subneural Operator (FSO) to output the whole image through multi-heads $H_{N}$. We denote frozen, reused, and trainable parameters in training at session 2. Note that each video representation is colored. In inference, we only need indices of session $s$ and frame $t$ and session mask (subnetwork).

### v4

@@ Abstract
Inspired by the Lottery Ticket Hypothesis (LTH), which highlights the existence of efficient subnetworks within larger, dense networks, a high-performing Winning Subnetwork (WSN) in terms of task performance under appropriate sparsity conditions is considered for various continual learning tasks. It leverages pre-existing weights from dense networks to achieve efficient learning in Task Incremental Learning (TIL) and Task-agnostic Incremental Learning (TaIL) scenarios. In Few-Shot Class Incremental Learning (FSCIL), a variation of WSN referred to as the Soft subnetwork (SoftNet) is designed to prevent overfitting when the data samples are scarce. Furthermore, the sparse reuse of WSN weights is considered for Video Incremental Learning (VIL). The use of Fourier Subneural Operator (FSO) within WSN is considered. It enables compact encoding of videos and identifies reusable subnetworks across varying bandwidths. We have integrated FSO into different architectural frameworks for continual learning, including VIL, TIL, and FSCIL. Our comprehensive experiments demonstrate FSO’s effectiveness, significantly improving task performance at various convolutional representational levels. Specifically, FSO enhances higher-layer performance in TIL and FSCIL and lower-layer performance in VIL.
@@
Continual Learning (CL), Task Incremental Learning (TIL),Task-agnostic Incremental Learning (TaIL), Video Incremental Learning (VIL), Few-shot Class Incremental Learning (FSCIL), Regularized Lottery Ticket Hypothesis (RLTH), Wining SubNetworks (WSN), Soft-Subnetwork (SoftNet), Fourier Subneural Operator (FSO)
@@ 1 Introduction
These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting or to alleviate overfitting. However, they are unsuited for sequential complex Video Incremental Learning (VIL) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we propose a novel sequential video compilation method to identify and utilize Lottery tickets (i.e., the weights of complex signals) in frequency space. To achieve this, we define Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. Given a backbone and FSO architecture, our method continuously learns to identify input-adaptive sub-modules in Fourier space and encode videos in each sequential training session. We apply FSO to various architectures accompanied by continual learning scenarios, such as Task Incremental Learning (TIL) and Task-agnostic Incremental Learning (TaIL), Video Incremental Learning (VIL), and Few-shot Class Incremental Learning (FSCIL), to demonstrate the effectiveness of FSO representations.
@@
  • We have applied the FSO to various architectures used by a variety of continual learning scenarios: Video Incremental Learning (VIL), Task Incremental Learning (TIL), Task-agnostic Incremental Learning (TaIL), and Few-Shot Class Incremental Learning (FSCIL). In our evaluations, the proposed FSO performs better than architecture-based continual learning models, such as WSN and SoftNet, in TIL, TaIL, VIL, and FSCIL, respectively, underscoring its exceptional representational power.
@@ 2 Related Works
Continual Learning [5, 1, 31, 32], also known as lifelong learning, is the challenge of learning a sequence of tasks continuously while utilizing and preserving previously learned knowledge to improve performance on new tasks. Four major approaches have been proposed to tackle the challenges of continual learning, such as catastrophic forgetting. One such approach is regularization-based methods [6, 7, 8, 9, 10], which aim to reduce catastrophic forgetting by imposing regularization constraints that inhibit changes to the weights or nodes associated with past tasks. Rehearsal-based approaches [11, 13, 14, 15, 33, 34, 16, 35, 36, 37, 38, 38, 39, 40, 41] store small data summaries to the past tasks and replay them during training to retain the acquired knowledge. Some methods in this line of work [42, 37] accommodate the generative model to construct the pseudo-rehearsals for previous tasks. Architecture-based approaches [17, 18, 19, 20, 21, 22] use the additional capacity to expand [43, 30], dynamic representation [44, 45] or isolate [2] model parameters, preserving learned knowledge and preventing forgetting. Both rehearsal and architecture-based methods have shown remarkable efficacy in suppressing catastrophic forgetting but require additional capacity for the task-adaptive parameters [20] or the replay buffers. Recently, Prompt-based learning, an emerging transfer learning technique in natural language processing (NLP), harnesses a fixed function of pre-trained Transformer models. This empowers the language model to receive additional instructions for enhancing its performance on downstream tasks. Notably, while L2P [46] stands out as the seminal work that bridges the gap between prompting and continual learning, DualPrompt [47] introduces an innovative approach to affixing complementary prompts to the pre-trained backbone, thereby enabling the acquisition of both task-invariant and task-specific instructions. Additionally, other notable contributions in this field encompass DyTox [48], S-Prompt [49], CODA-P [50], ConStruct-VL [51], ST-Prompt [52], LGCL [53], PGP [54]. All previous learning methods depend mainly on continual representations in real space. However, the central focus of this study is to pinpoint the most optimal winning ticket representations for convolutional operators in four continual learning scenarios in Fourier space.
Architecture-based Continual Learning. Artificial network architecture is designed to enable the training of deeper networks. ResNets [55], as a fundamental backbone network with a convolutional operator that shares parameters to obtain image representations in latent spaces, have been widely used in various research fields, such as image classification [56, 57], object detection [58, 59], semantic segmentation [60, 61], image captioning [62, 63], image generation [64, 65], and architecture-based continual learning [24, 17, 18, 66]. A recent CL method, LL-Tickets [66], shows a sparse subnetwork called lifelong tickets that performs well on all tasks during continual learning. However, LL-Tickets require external data to maximize knowledge distillation with learned models for prior tasks, and the ticket expansion process involves retraining and pruning steps. WSN [21] was an improved method that jointly learns the model and task-adaptive subnetwork associated with each task in Task Incremental Learning (TIL). Also, in Few-Shot Class Incremental Learning (FSCIL), Soft-Subnetworks (SoftNet) [22] was proposed as another variant of WSN, consisting of major sub-networks (winning tickets) to obtain base session knowledge and minor sub-networks to alleviate overfitting few samples for new sessions. However, these methods adhere to the traditional ResNet architecture. As the layers increase in depth, the image representations become more abstract, which can result in the loss of the global structure of the input images. These drawbacks can negatively impact the effectiveness of image representation and the continuity of the representational power. To overcome these issues, we introduce a new convolutional operator referred to as the Fourier Subneural Operator (FSO), which transfers the global image representation obtained from lower layers to higher layers through Residual Blocks. This approach helps maintain the global image representation in Artificial Neural Networks at higher layers.
@@
Neural Implicit Representation (NIR) [67] are neural network architectures for parameterizing continuous, differentiable signals. Based on coordinate information, they provide a way to represent complex, high-dimensional data with a small set of learnable parameters that can be used for various tasks such as image reconstruction [68, 69], shape regression [70, 71], and 3D view synthesis  [72, 73]. Instead of using coordinate-based methods, NeRV [74] proposes an image-wise implicit representation that takes frame indices as inputs, enabling faster and more accurate video compression than coordinate methods. NeRV has inspired further improvements in video regression by CNeRV [75], DNeRV [76], E-NeRV [77], and NIRVANA [78], and HNeRV [79]. A few recent works have explored video continual learning (VCL) scenarios for the NIR. To tackle non-physical environments, Continual Predictive Learning (CPL) [80] learns a mixture world model via predictive experience replay and performs test-time adaptation using non-parametric task inference. PIVOT [81] leverages the past knowledge present in pre-trained models from the image domain to reduce the number of trainable parameters and mitigate forgetting. CPL needs memory to replay, while PIVOT needs pre-training and fine-tuning steps. In contrast, along with the conventional progressive training techniques [2, 82], considering the advantages of forget-free convergence speed, we set WSN as baselines, which utilizes the Lottery Ticket Hypothesis (LTH) to identify an adaptive substructure within the dense networks that are tailored to the specific video input index. However, WSN is inappropriate for sequential complex video compilation since it reuses a few adaptive but sparse learnable parameters. To overcome the weakness of WSN, We proposed a novel Fourier Subneural Operator (FSO) [83] for representing complex video in Fourier space [84, 85, 86, 87]. We have expanded the FSO of Fourier representations to encompass a variety of continual learning architectures and scenarios to validate its effectivness.
@@ 3 Winning SubNetworks in Fourier Space
This section presents our pruning-based continual learning methods, Winning SubNetworks (WSN, see Figure 2[21] and introduces the Fourier Subneural Operator (FSO, see Figure 3 and Figure 4) for better video representations. Then, we depict how we apply FSO to various architectures used in four continual learning scenarios.
@@
Figure 2: An illustration of Winning SubNetworks (WSN): (a) The top-cweights $\hat{\bm{\theta}}_{s-1}$ at prior task are obtained, (b) In the forward pass of a new task, WSN selects the top-cand reuses weights selected from prior tasks, (c) In the backward pass, WSN updates only non-used weights( ) while freezing reused weights( ), and (d) after several iterations of (b) and (c), we acquire again the top-cweights $\hat{\bm{\theta}}_{s}$ including subsets of reused weights (green) for the new task.
3.1 WSN & Fourier Subneural Operator (FSO)
The WSN searches for the task-adaptive winning tickets and updates only the weights not trained on the previous tasks, as shown in Figure 2. After training on each session, the subnetwork parameters of the model are frozen to ensure that the proposed method is inherently immune to catastrophic forgetting. Moreover, we illustrate Soft-Winning SubNetworks (SoftNet) [22], proposed to address the issues of forgetting previous sessions and overfitting a few samples of new sessions. These conventional architecture-based methods, i.e., WSN and SoftNet, offer solutions to prevent forgetting. However, they are unsuited for sequential complex Video Incremental Learning (VIL see Figure 3) as they reuse a few or all adaptive parameters without finely discretized operations. To enhance neural representation incrementally on complex sequential videos, we introduce Fourier Subneural Operator (FSO), which breaks down a neural implicit representation into its sine and cosine components (real and imaginary parts) and then selectively identifies the most effective Lottery tickets for representing complex periodic signals. In practice, given a backbone and FSO architecture, our method continuously learns to identify input-adaptive subnetwork modules and encode each new video into the corresponding module during sequential training sessions. We extend Fourier representations to various continual learning scenarios, such as TIL, TaIL, VIL, and FSCIL, to demonstrate its effectiveness.
Problem Statement. Consider a supervised learning setup where $S$ sessions or tasks arrive to a learner sequentially. We denote that $\mathcal{D}_{s}=\{\bm{x}_{i,s},y_{i,s}\}_{i=1}^{n_{s}}$ is the dataset of session $s$, composed of $n_{s}$ pairs of raw instances and corresponding labels. We assume a neural network $f(\cdot;\bm{\theta})$, parameterized by the model weights $\bm{\theta}$ and standard continual learning scenario aims to learn a sequence of sessions by solving the following optimization procedure at each step $s$:
$\bm{\theta}^{\ast}=\operatorname*{minimize}_{\bm{\theta}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}(f(\bm{x}_{i,s};\bm{\theta}),y_{i,s}),$ (1)
@@ 3.1 WSN & Fourier Subneural Operator (FSO)
Continual learners frequently use over-parameterized deep neural networks (dense network) to ensure enough capacity for learning future tasks. This approach often leads to the discovery of subnetworks that perform as well as, or better than, the dense network. Given the neural network parameters $\bm{\theta}$, the binary mask $\bm{m}^{*}_{s}$ that describes the optimal subnetwork for session $s$ such that $|\bm{m}^{*}_{s}|$ is less than the model capacity $c$, is defined as:
$\begin{split}\bm{m}^{*}_{s}&=\underset{\bm{m}_{s}\in\{0,1\}^{|\bm{\theta}|}}{\operatorname*{minimize}}\frac{1}{n_{s}}\sum^{n_{s}}_{i=1}\mathcal{L}\big{(}f(\bm{x}_{i,s};\bm{\theta}\odot\bm{m}_{s}),y_{i,s}\big{)}-\mathcal{J}\\ &\quad\quad\quad\quad\quad\quad\text{subject to\leavevmode\nobreak\ }|\bm{m}^{*}_{s}|\leq c\ll|\bm{\theta}|,\end{split}$ (2)
@@ 3.2 Fourier Subnueral Operator (FSO)
$\left(\mathcal{K}(\phi)\tilde{\bm{v}}_{t}^{s}\right)(\bm{e}_{s,t})=\mathcal{F}^{-1}(R_{\bm{\phi}}\cdot(\mathcal{F}\tilde{\bm{v}}_{t}^{s}))(\bm{e}_{s,t}),$ (3)
where $\tilde{\bm{v}}_{t}^{s}$ is a hidden representation; $R_{\bm{\phi}}$ is the Fourier transform of a periodic subnetwork function which is parameterized by its subnetwork’s parameters of real $(\bm{\theta}^{real}\odot\bm{m}_{s}^{real})$ and imaginary $(\bm{\theta}^{imag}\odot\bm{m}_{s}^{imag})$. We thus parameterize $R_{\bm{\phi}}$ separately as complex-valued tensors of real and imaginary $\bm{\phi}_{FSO}\in\{\bm{\theta}^{real},\bm{\theta}^{imag}\}$. One key aspect of the FSO is that its parameters grow with the depth of the layer and the input/output size. However, through careful layer-wise inspection and adjustments for sparsity, we can find a balance that allows the FSO to describe neural implicit representations efficiently. In the experimental section, we will showcase the most efficient FSO structure and its performance. Figure 3 shows one possible structure of a single FSO. We describe the optimization in the following section.
Figure 3: Forget-free Neural Implicit Representation with Fourier Subneural Operator (FSO) for Video Incremental Learning: Image-wise neural implicit representation taking frame and video (session $s$) indices as input and using a sparse Stems + NeRV Blocks with Fourier Subneural Operator (FSO) to output the whole image through multi-heads $H_{N}$ where $\tilde{\bm{v}}_{s}^{t}$ is a hidden representation. We denote frozen, reused, and trainable parameters in training at session 2. Note that each video representation is colored. In inference, we only need indices of session $s$ and frame $t$ and session mask (subnetwork).